# B-cos LM: Efficiently Transforming Pre-trained Language Models for Improved Explainability

### v3

@@ Abstract
Post-hoc explanation methods for black-box models often struggle with faithfulness and human interpretability due to the lack of explainability in current neural architectures. Meanwhile, B-cos networks have been introduced to improve model explainability by proposing an architecture that removes bias terms and promotes input-weight alignment. Although B-cos networks have shown success in building explainable systems, their application has so far been limited to computer vision models and their associated training pipelines. In this work, we introduce B-cos LMs, i.e., B-cos language models (LMs) empowered for natural language processing (NLP) tasks. Our approach directly transforms pre-trained language models into B-cos LMs by combining B-cos conversion and task fine-tuning, improving efficiency compared to previous methods. Our automatic and human evaluation results demonstrate that B-cos LMs produce more faithful and human interpretable explanations than post-hoc methods, while maintaining task performance comparable to conventional fine-tuning. Our in-depth analysis explores how B-cos LMs differ from conventionally fine-tuned models in their learning processes and explanation patterns. Finally, we present a first exploration of transforming decoder-only models to B-cos LMs for generation tasks.
@@ 1 Introduction
Pre-trained language models (PLMs) such as BERT (Devlin et al., 2019) and GPT (Radford et al., 2019; Brown et al., 2020; OpenAI, 2023) have significantly advanced performance across a plethora of NLP tasks (Wang et al., 2018; Gao et al., 2023). However, their complex architectures and black-box nature make understanding their behavior a persistent challenge (Bommasani et al., 2021). To address this, research has increasingly focused on understanding model predictions in various natural language understanding and generation tasks using different forms of explanations, such as input-based explanations (Feng et al., 2024; Wei Jie et al., 2024; Jiang et al., 2024; Madsen et al., 2024; Yin & Neubig, 2022; Deiseroth et al., 2023), natural language explanations (Ramnath et al., 2024; Wang et al., 2025), and concept-based explanations (Yu et al., 2024; Raman et al., 2024). Among others, input-based explanations, often referred to as rationales, aim to reveal how specific inputs influence a model’s prediction (Arras et al., 2019; Atanasova et al., 2020; Lyu et al., 2024). In this work, we focus on input-based explanations, as they offer the most direct insight into model behavior and are often mandated by laws, such as the EU Artificial Intelligence Act.
@@
Most input-based explanation methods for neural models are post-hoc, meaning that they attempt to explain a model’s behavior only after it has been trained and deployed (Sundararajan et al., 2017; Ribeiro et al., 2016). While these methods are widely used and easy to apply, they have been shown to produce unfaithful explanations that do not accurately reflect the model’s actual reasoning process (Kindermans et al., 2019; Slack et al., 2020; Pruthi et al., 2020). They also struggle with human interpretability, making it difficult for users to understand the model’s reasoning (Smilkov et al., 2017; Ismail et al., 2021). Prior research suggests that these limitations stem from a lack of inherent explainability in current models, that is, the model’s ability to generate faithful and interpretable explanations by design (Kindermans et al., 2018; Alvarez Melis & Jaakkola, 2018; Rudin, 2019). As a result, improving model explainability is crucial for producing explanations that are both reliable and useful to users. [^1] Considering the evolving definition of these terms in past literature, we provide a detailed definition in Appendix A. Figure 1 provides examples illustrating this issue.
To overcome these limitations, we introduce B-cos LM, a dynamic linear model that learns the most task-relevant patterns through increased input-weight alignment pressure. Building upon B-cos networks that were first introduced by Böhle et al. (2022) for computer vision, we ensure the explainability of B-cos LMs through mathematically grounded architectural and computational adaptations, with specialized architectural modifications and training pipelines tailored for NLP tasks.
We conduct comprehensive empirical experiments using encoder-only models on classification tasks. Our focus on classification is motivated by its prevalence in high-stakes applications, such as loan approvals, hiring decisions, and hate speech detection, where explainability is crucial and often legally mandated. Encoder-only models have also seen renewed interest in the research community (Warner et al., 2024; Breton et al., 2025; Chaffin, 2025), and they remain the standard architecture for text classification and continue to perform competitively compared to large language models (LLMs) (Zhao et al., 2024). Beyond that, we also explore applying B-cos LMs to decoder-only models for generation tasks and show that B-cos LMs can be extended to a variety of tasks and the latest model architectures. Our contributions are as follows:
@@ 2 Related Work
Various methods have been proposed to provide post-hoc explanations for neural model predictions (Atanasova et al., 2020). These methods can be broadly categorized based on how they generate explanations: gradient-based (Simonyan et al., 2014; Kindermans et al., 2016; Sundararajan et al., 2017; Enguehard, 2023), propagation-based (Bach et al., 2015; Shrikumar et al., 2017; Springenberg et al., 2015; Ferrando et al., 2023; Modarressi et al., 2022; 2023), and perturbation-based methods (Li et al., 2016; Ribeiro et al., 2016; Lundberg & Lee, 2017; Deiseroth et al., 2023). Besides, the attention mechanism (Bahdanau et al., 2015) is often viewed as an explanation, particularly in transformer-based models (Vaswani et al., 2017). While most existing work focuses on understanding model predictions in classification settings, recent efforts have also aimed to explain model behavior in generation tasks, including sentence completion (Yin & Neubig, 2022; Ferrando et al., 2023), question answering (Enouen et al., 2024), and summarization (Cohen-Wang et al., 2024).
Although post-hoc methods have been widely used, numerous studies have shown that they lack faithfulness, often failing to capture the true decision-making process of the model (Kindermans et al., 2019; Jain & Wallace, 2019; Slack et al., 2020; Pruthi et al., 2020). Furthermore, they are noisy and may select irrelevant information leading to explanations that cannot be interpreted by humans (Smilkov et al., 2017; Ismail et al., 2021).
@@
Prior research suggests that the lack of faithfulness and human interpretability in post-hoc explanations arises from the fundamental lack of explainability in modern neural models, which are typically optimized solely for task performance (Kindermans et al., 2018; Rudin, 2019; Atanasova et al., 2022). In response, various efforts have been made to enhance model explainability. Some works have introduced constraints that improve specific explanation properties, such as faithfulness (Tutek & Šnajder, 2022; Moradi et al., 2020; 2021; Barkan et al., 2024), consistency (Atanasova et al., 2022), locality (Alvarez Melis & Jaakkola, 2018), and plausibility (Ismail et al., 2021). However, as these constraints are typically imposed as regularizers, their effectiveness in improving explanation quality is not guaranteed (Pruthi et al., 2020). Others have proposed self-explanatory model architectures such as rationale-based models that utilize an “explain-then-predict” pipeline, where one module selects rationales for another to make predictions based on them (Lei et al., 2016). Although seemingly transparent, both components rely on neural networks, making the rationale extraction and utilization processes opaque (Zheng et al., 2022; Jacovi & Goldberg, 2021). Besides, such models may face optimization challenges that limit their practicality in real-world tasks (Lyu et al., 2024).
To tackle these shortcomings, Böhle et al. (2022) proposed B-cos networks. Unlike methods that impose external constraints, B-cos networks improve explainability through mathematically grounded architectural and computational adaptations. Moreover, these adaptations are designed as drop-in replacements for conventional model components, making B-cos networks easy to train with minimal performance loss. Most recently, Arya et al. (2024) explored B-cosification techniques to convert existing models into B-cos models, which reduces the training costs of adopting B-cos architectures.
@@ 3.1 B-cos Networks
Complex neural networks can be interpreted as generalized linear models (Nair & Hinton, 2010; Alvarez Melis & Jaakkola, 2018; Srinivas & Fleuret, 2019). For each input $\mathbf{x}$, the network effectively applies a linear transformation: $\mathbf{f(x)=W(x)x+b(x)}$, where both the weight $\mathbf{W(x)}$ and bias $\mathbf{b(x)}$ depend on $\mathbf{x}$. Given that many activation functions are (approximately) piecewise linear, the overall network can be viewed as (approximately) piecewise affine (Alvarez Melis & Jaakkola, 2018). Earlier work refers to such models as dynamic linear models (Böhle et al., 2021; 2022), highlighting the fact that the weight and bias terms dynamically change according to $\mathbf{x}$.
Under this dynamic linear perspective, the linear mapping $\mathbf{W(x)}$ can be seen as attributing model predictions to individual input features, and $\mathbf{W(x_{i})x_{i}}$ can be seen as the contribution of feature $\mathbf{x_{i}}$ to the model prediction. However, two challenges hinder the direct use of this interpretation. First, $\mathbf{W(x)}$ alone provides an incomplete and unfaithful model summary since $\mathbf{f(x)\neq W(x)x}$ due to the presence of the bias term $\mathbf{b(x)}$, and incorporating $\mathbf{b(x)}$ into explanations is highly non-trivial (Wang et al., 2019). Second, $\mathbf{W(x_{i})x_{i}}$ is often difficult for humans to interpret, as $\mathbf{W(x)}$ does not necessarily align only with task-relevant input patterns (Smilkov et al., 2017) and therefore yields noisy and irrelevant explanations. Figure 1 illustrates these challenges. To address these issues, Böhle et al. (2022) introduced B-cos networks by replacing the conventional linear transformation:
@@
While early B-cos models were trained from scratch, Arya et al. (2024) recently introduced B-cosification, an efficient method to obtain B-cos models. This approach first modifies conventional models with task capacities to adopt the B-cos architecture, followed by fine-tuning on downstream datasets for B-cos conversion. B-cosified models generate explanations as faithful and interpretable as B-cos models trained from scratch but at a much lower training cost. However, directly applying B-cosification to LMs is non-trivial and inefficient due to the significant differences in model architectures and training pipelines.
@@ 3.2 B-cosification for LMs
Property Conventional Fine-tuning B-cosification for vision (Arya et al., 2024) B-cos LM (ours)
@@
Pred. Head Activations tanh n/a [^2] Arya et al. (2024) used a single linear layer on top of CLIP so the prediction head activation is not applicable in their setup. identity
@@
Table 1: Comparison between conventional fine-tuning, B-cosification for computer vision models and B-cosification for language models (B-cos LM). Conventional fine-tuning and B-cosification for vision follow the configuration of BERT for sequence classification and CLIP (Radford et al., 2021), respectively (cf. § 3 for details).
@@ 3.2.1 B-cos Adaptations
For completeness and faithfulness of explanations, we follow Arya et al. (2024) and remove all bias terms in models, including those in the affine transformations of layer normalization and attention blocks. Additionally, a prediction head is typically added on top of the transformer before fine-tuning for downstream tasks in the NLP pipeline. This head often includes activation functions that are not (approximately) piecewise linear, such as sigmoid and tanh. To accommodate the unique architecture of LMs, we remove all activation functions in the prediction heads, as they generate explanations that are not locally difference-bounded (Alvarez Melis & Jaakkola, 2018) and introduce numerical instability during explanation generation. Our experiments show that the added non-linearity from B>1 could compensate for this removal.
@@ 3.2.2 Fine-tuning
The B-cos adaptations above modify the architecture and computation of models, requiring fine-tuning to restore their capabilities and adapt to alignment pressure. Following the “pre-train then fine-tune” paradigm, which is frequently utilized in NLP tasks, we directly transform PLMs to B-cos LMs, rather than adapting task-specific models as done in previous work (Arya et al., 2024). This fundamental difference in the training pipeline adds complexity to B-cosification for LMs, as the objective involves both B-cos conversion and task fine-tuning. While there are multiple ways to conjoin these two steps (cf. § 5), we find that the most efficient way is to combine them by first applying B-cos adaptations to a PLM and then fine-tuning it on a downstream task. Following Böhle et al. (2022), we use the binary cross-entropy (BCE) loss instead of the conventional cross-entropy loss, as it explicitly maximizes the absolute target logits and strengthens the alignment pressure. We provide an extensive comparison of different B-cosification setups in § 5.
@@ 4.1 Experimental Setup
Our experiments use three datasets: AG News (topic classification, Zhang et al., 2015), IMDB (sentiment analysis, Maas et al., 2011), and HateXplain (hate speech detection, Mathew et al., 2021). BERT (Devlin et al., 2019), DistilBERT (Sanh et al., 2019), and RoBERTa (Liu et al., 2019) are used as the basis for conventional fine-tuning and for obtaining B-cos LMs. We set B=1.25 for IMDB and B=1.5 for AG News and HateXplain datasets.
@@
We compare B-cos explanations against a diverse set of post-hoc explanation methods: Attention (Bahdanau et al., 2015), InputXGradient (IxG, Kindermans et al., 2016), Sequential Integrated Gradients (SIG, Enguehard, 2023), DecompX (Modarressi et al., 2023), Shapley Value Sampling (ShapSampl, Strumbelj & Kononenko, 2010), and LIME (Ribeiro et al., 2016). We also apply these methods to a model trained with Saloss (Chrysostomou & Aletras, 2021), which incorporates additional faithfulness regularization. This setup enables a direct comparison between B-cos LMs and models specifically optimized for explainability. For embedding-level explanation methods, we aggregate attributions by summing across all embedding dimensions.
@@
For a more comprehensive evaluation, we employ two different methods to assess faithfulness. First, we report two perturbation-based metrics (DeYoung et al., 2020):
@@
To avoid arbitrary choices of $k$, we compute Comp and Suff for multiple values (k=10,20,,90k=10,20,...,90) and summarize them using the Area Over the Perturbation Curve (AOPC, DeYoung et al., 2020).
In addition, we introduce a new faithfulness metric called Sequence Pointing Game (SeqPG), inspired by the grid pointing game in vision tasks (Böhle et al., 2021):
@@
Compared to perturbation-based metrics, SeqPG does not rely on perturbations and thus avoids the potential distortions introduced by token masking. When constructing SeqPG examples, we truncate each segment to a fixed length and randomize segment order to control for length and position effects. We generate synthetic examples using correctly and most confidently classified test instances. SeqPG can be seen as a standardized version of hybrid document evaluation (Poerner et al., 2018). We provide an example of SeqPG in Figure 7 and more details in Appendix F.

### v4

@@ Abstract
Post-hoc explanation methods for black-box models often struggle with faithfulness and human interpretability due to the lack of explainability in current neural architectures. Meanwhile, B-cos networks have been introduced to improve model explainability by proposing an architecture that removes bias terms and promotes input-weight alignment. Although B-cos networks have shown success in building explainable systems, their application has so far been limited to computer vision models and their associated training pipelines. In this work, we introduce B-cos LMs, i.e., B-cos Language Models (LMs) empowered for natural language processing (NLP) tasks. Our approach directly transforms pre-trained language models into B-cos LMs by combining B-cos conversion and task fine-tuning, improving efficiency compared to previous methods. Automatic and human evaluation results demonstrate that B-cos LMs produce more faithful and human interpretable explanations than post-hoc methods, while maintaining task performance comparable to conventional fine-tuning. Our in-depth analysis explores how B-cos LMs differ from conventionally fine-tuned models in their learning processes and explanation patterns. Finally, we present a first exploration of transforming decoder-only models to B-cos LMs for generation tasks. Our code is available at https://github.com/Ewanwong/bcos_lm.
@@ 1 Introduction
Pre-trained language models (PLMs) such as BERT (Devlin-2019-BERT) and GPT (Radford-2019-GPT; NEURIPS2020_1457c0d6; DBLP:journals/corr/abs-2303-08774) have significantly advanced performance across a plethora of NLP tasks (wang-etal-2018-glue; Eval-2023-Harness). However, their complex architectures and black-box nature make understanding their behavior a persistent challenge (Bommasani-2021-FoundationModels). To address this, research has increasingly focused on understanding model predictions in various natural language understanding and generation tasks using different forms of explanations, such as input-based explanations (zijian-etal-2024-unveiling; wei-jie-etal-2024-plausible; jiang-etal-2024-mare; madsen-etal-2024-self; yin-neubig-2022-interpreting; deiseroth-etal-2023-atman; abbasi-etal-2025-normxlogit), natural language explanations (sahana-etal-2024-tailoring; wang-etal-2025-cross), and concept-based explanations (yu-etal-2024-latent; raman-etal-2024-understanding). Among others, input-based explanations, often referred to as rationales, aim to reveal how specific inputs influence a model’s prediction (arras-etal-2019-evaluating; atanasova-etal-2020-diagnostic; lyu-etal-2024-towards). In this work, we focus on input-based explanations, as they offer the most direct insight into model behavior and are often mandated by laws, such as the EU Artificial Intelligence Act.
@@
Most input-based explanation methods for neural models are post-hoc, meaning that they attempt to explain a model’s behavior only after it has been trained and deployed (lyu-etal-2024-towards). While these methods are widely used and easy to apply, they have been shown to produce unfaithful explanations that do not accurately reflect the model’s actual reasoning process (Kindermans-2019-Reliability; Slack-2020-FoolingLimeSHAP; pruthi-etal-2020-learning; ye-etal-2025-input). They also struggle with human interpretability, making it difficult for users to understand the model’s reasoning (Smilkov-2017-SmoothGrad; NEURIPS2021_e0cd3f16). Prior research suggests that these limitations stem from a lack of inherent explainability in current models, that is, the model’s ability to generate faithful and interpretable explanations by design (Kindermans-2018-PatternNet; NEURIPS2018_3e9f0fc9; rudin2019stop). As a result, improving model explainability is crucial for producing explanations that are both reliable and useful to users. [^1] Considering the evolving definition of these terms in past literature, we provide a detailed definition in Appendix A. Figure 1 provides examples illustrating this issue.
To overcome these limitations, we introduce B-cos LM, a dynamic linear model that learns the most task-relevant patterns through increased input-weight alignment pressure. Building upon B-cos networks that were first introduced by Bohle_2022_CVPR for computer vision, we ensure the explainability of B-cos LMs through mathematically grounded architectural and computational adaptations, with specialized architectural modifications and training pipelines tailored for NLP tasks.
We conduct comprehensive empirical experiments using encoder-only models on classification tasks. Our focus on classification is motivated by its prevalence in high-stakes applications, such as loan approvals, hiring decisions, and hate speech detection, where explainability is crucial and often legally mandated. Encoder-only models have also seen renewed interest in the research community (modernbert; neobert; Reason-ModernColBERT), and they remain the standard architecture for text classification tasks and continue to perform competitively compared to large language models (LLMs) (hang-etal-2024-advancing). Beyond that, we also explore applying B-cos LMs to decoder-only models for generation tasks and show that B-cos LMs can be extended to a variety of tasks and the latest model architectures. Our contributions are as follows:
@@ 2 Related Work
Various methods have been proposed to provide post-hoc explanations for neural model predictions (atanasova-etal-2020-diagnostic). These methods can be broadly categorized based on how they generate explanations: gradient-based (DBLP:journals/corr/SimonyanVZ13; DBLP:journals/corr/KindermansSMD16; Sundararajan-2017-IntegratedGrad; enguehard-2023-sequential), propagation-based (bach2015pixel; shrikumar2017learning; DBLP:journals/corr/SpringenbergDBR14; ferrando-etal-2023-explaining; modarressi-etal-2022-globenc; modarressi-etal-2023-decompx), and perturbation-based methods (DBLP:journals/corr/LiMJ16a; Ribeiro-2016-LIME; Lundberg-2017-SHAP; deiseroth-etal-2023-atman). Besides, the attention mechanism (DBLP:journals/corr/BahdanauCB14) is often viewed as an explanation, particularly in transformer-based models (DBLP:conf/nips/VaswaniSPUJGKP17). While most existing work focuses on understanding model predictions in classification settings, recent efforts have also aimed to explain model behavior in generation tasks, including sentence completion (yin-neubig-2022-interpreting; ferrando-etal-2023-explaining), question answering (enouen-etal-2024-textgenshap), and summarization (cohen-etal-2024-contextcite).
Although post-hoc methods have been widely used, numerous studies have shown that they lack faithfulness, often failing to capture the true decision-making process of the model (Kindermans-2019-Reliability; jain-wallace-2019-attention; Slack-2020-FoolingLimeSHAP; pruthi-etal-2020-learning; ye-etal-2025-input). Furthermore, they are noisy and may select irrelevant information, leading to explanations that cannot be interpreted by humans (Smilkov-2017-SmoothGrad; NEURIPS2021_e0cd3f16).
@@
Prior research suggests that the lack of faithfulness and human interpretability in post-hoc explanations arises from the fundamental lack of explainability in modern neural models, which are typically optimized solely for task performance (Kindermans-2018-PatternNet; rudin2019stop; atanasova2022diagnostics). In response, various efforts have been made to enhance model explainability. Some works have introduced constraints that improve specific explanation properties, such as faithfulness (tutek2022toward; moradi-etal-2020-training; moradi-etal-2021-measuring; barkan-etal-2024-llm), consistency (atanasova2022diagnostics), locality (NEURIPS2018_3e9f0fc9), and plausibility (NEURIPS2021_e0cd3f16). However, as these constraints are typically imposed as regularizers, their effectiveness in improving explanation quality is not guaranteed (pruthi-etal-2020-learning). Others have proposed self-explanatory model architectures such as rationale-based models that utilize an “explain-then-predict” pipeline, where one module selects rationales for another to make predictions based on them (lei-etal-2016-rationalizing). Although seemingly transparent, both components rely on neural networks, making the rationale extraction and utilization processes opaque (zheng-etal-2022-irrationality; jacovi-goldberg-2021-aligning). Besides, such models may face optimization challenges that limit their practicality in real-world tasks (lyu-etal-2024-towards).
To tackle these shortcomings, Bohle_2022_CVPR proposed B-cos networks. Unlike methods that impose external constraints, B-cos networks improve explainability through mathematically grounded architectural and computational adaptations. Moreover, these adaptations are designed as drop-in replacements for conventional model components, making B-cos networks easy to train with minimal performance loss. Most recently, arya24bcosification explored B-cosification techniques to convert existing models into B-cos models, which reduces the training costs of adopting B-cos architectures.
@@ 3.1 B-cos Networks
Complex neural networks can be interpreted as generalized linear models (DBLP:conf/icml/NairH10; NEURIPS2018_3e9f0fc9; NEURIPS2019_80537a94). For each input $\mathbf{x}$, the network effectively applies a linear transformation: $\mathbf{f(x)=W(x)x+b(x)}$, where both the weight $\mathbf{W(x)}$ and bias $\mathbf{b(x)}$ depend on $\mathbf{x}$. Given that many activation functions are (approximately) piecewise linear, the overall network can be viewed as (approximately) piecewise affine (NEURIPS2018_3e9f0fc9). Earlier work refers to such models as dynamic linear models (bohle2021convolutional; Bohle_2022_CVPR), highlighting the fact that the weight and bias terms dynamically change according to $\mathbf{x}$.
Under this dynamic linear perspective, the linear mapping $\mathbf{W(x)}$ can be seen as attributing model predictions to individual input features, and $\mathbf{W(x_{i})x_{i}}$ can be seen as the contribution of feature $\mathbf{x_{i}}$ to the model prediction. However, two challenges hinder the direct use of this interpretation. First, $\mathbf{W(x)}$ alone provides an incomplete and unfaithful model summary, since $\mathbf{f(x)\neq W(x)x}$ due to the presence of the bias term $\mathbf{b(x)}$, and incorporating $\mathbf{b(x)}$ into explanations is highly non-trivial (pmlr-v97-wang19p). Second, $\mathbf{W(x_{i})x_{i}}$ is often difficult for humans to interpret, as $\mathbf{W(x)}$ does not necessarily align only with task-relevant input patterns (Smilkov-2017-SmoothGrad) and therefore yields noisy and irrelevant explanations. Figure 1 illustrates these challenges. To address these issues, Bohle_2022_CVPR introduced B-cos networks by replacing the conventional linear transformation:
@@
While early B-cos models were trained from scratch, arya24bcosification recently introduced B-cosification, an efficient method to obtain B-cos models. This approach first modifies conventional models with task capacities to adopt the B-cos architecture, followed by fine-tuning on downstream datasets for B-cos conversion. B-cosified models generate explanations as faithful and interpretable as B-cos models trained from scratch but at a much lower training cost. However, directly applying B-cosification to LMs is non-trivial and inefficient due to the significant differences in model architectures and training pipelines.
@@ 3.2 B-cosification for LMs
Property Conventional Fine-tuning B-cosification for vision (arya24bcosification) B-cos LM (ours)
@@
Pred. Head Activations tanh n/a222arya24bcosification used a single linear layer on top of CLIP so the prediction head activation is not applicable in their setup. identity
@@
Table 1: Comparison between conventional fine-tuning, B-cosification for vision models and B-cosification for language models (B-cos LM). Conventional fine-tuning and B-cosification for vision follow the configuration of BERT for sequence classification and CLIP (radford2021learning), respectively (cf. § 3 for details).
@@ 3.2.1 B-cos Adaptations
For completeness and faithfulness of explanations, we follow arya24bcosification and remove all bias terms in models, including those in the affine transformations of layer normalization and attention blocks. Additionally, a prediction head is typically added on top of transformers before fine-tuning for downstream tasks in the NLP pipeline. This head often includes activation functions that are not (approximately) piecewise linear, such as sigmoid and tanh. To accommodate the unique architecture of LMs, we remove all activation functions in the prediction heads, as they generate explanations that are not locally difference-bounded (NEURIPS2018_3e9f0fc9) and introduce numerical instability during explanation generation. Our experiments show that the added non-linearity from B>1 could compensate for this removal.
@@ 3.2.2 Fine-tuning
The B-cos adaptations above modify the architecture and computation of models, requiring fine-tuning to restore their capabilities and adapt to alignment pressure. Following the “pre-train then fine-tune” paradigm, which is frequently utilized in NLP tasks, we directly transform PLMs to B-cos LMs, rather than adapting task-specific models as done in previous work (arya24bcosification). This fundamental difference in the training pipeline adds complexity to B-cosification for LMs, as the objective involves both B-cos conversion and task fine-tuning. While there are multiple ways to conjoin these two steps (cf. § 5), we find that the most efficient way is to combine them by first applying B-cos adaptations to a PLM and then fine-tuning it on a downstream task. Following Bohle_2022_CVPR, we use the binary cross-entropy (BCE) loss instead of the conventional cross-entropy loss, as it explicitly maximizes the absolute target logits and strengthens the alignment pressure. We provide an extensive comparison of different B-cosification setups in § 5.
@@ 4.1 Experimental Setup
Our experiments use three datasets: AG News (topic classification, NIPS2015_250cf8b5), IMDB (sentiment analysis, maas-etal-2011-learning), and HateXplain (hate speech detection, mathew2021HateXplain). BERT (Devlin-2019-BERT), DistilBERT (DBLP:journals/corr/abs-1910-01108), and RoBERTa (DBLP:journals/corr/abs-1907-11692) are used as the basis for conventional fine-tuning and for obtaining B-cos LMs. We set B=1.25 for IMDB and B=1.5 for AG News and HateXplain datasets.
@@
We compare B-cos explanations against a diverse set of post-hoc explanation methods: Attention (DBLP:journals/corr/BahdanauCB14), InputXGradient (IxG, DBLP:journals/corr/KindermansSMD16), Sequential Integrated Gradients (SIG, enguehard-2023-sequential), DecompX (modarressi-etal-2023-decompx), Shapley Value Sampling (ShapSampl, strumbelj2010efficient), and LIME (Ribeiro-2016-LIME). We also apply these methods to a model trained with Saloss (chrysostomou-aletras-2021-enjoy), which introduces faithfulness regularization. This setup enables a direct comparison between B-cos LMs and models specifically optimized for explainability. For embedding-level explanation methods, we aggregate attributions by summing across all embedding dimensions.
@@
For a more comprehensive evaluation, we employ two different methods to assess faithfulness. First, we report two perturbation-based metrics (deyoung-etal-2020-eraser):
@@
To avoid arbitrary choices of $k$, we compute Comp and Suff for multiple values (k=10,20,,90k=10,20,...,90) and summarize them using the Area Over the Perturbation Curve (AOPC, deyoung-etal-2020-eraser).
In addition, we introduce a new faithfulness metric called Sequence Pointing Game (SeqPG), inspired by the grid pointing game in vision tasks (bohle2021convolutional):
@@
Compared to perturbation-based metrics, SeqPG does not rely on perturbations and thus avoids the potential distortions introduced by token masking. When constructing SeqPG examples, we truncate each segment to a fixed length and randomize segment order to control for length and position effects. We generate synthetic examples using correctly and most confidently classified test instances. SeqPG can be seen as a standardized version of hybrid document evaluation (poerner-etal-2018-evaluating). We provide an example of SeqPG in Figure 7 and more details in Appendix F.